# Efficient Transfer Learning for Video-language Foundation Models

### v1

@@ Abstract
[…]
@@ 1 Introduction
Multi-modal foundation models (e.g., CLIP [28], ViCLIP [36], BLIP2 [21]) that are trained on large-scale datasets of paired multi-modal data have shown exceptional generalization capabilities. The semantic visual concepts empowered by large-scale multi-modal alignment can be widely applied to various downstream tasks, such as few-shot image/video classification [19, 3, 43, 29], open-vocabulary detection [12, 39], and segmentation [23]. In this paper, we focus on transferring the pre-trained multi-modal foundation models model (i.e., CLIP [28] and ViCLIP [36]) to the domain of video action recognition, emphasizing its crucial role in driving advancements in this field.
@@
The key to the adaptation process of foundational multi-modal models lies in injecting domain-specific expertise. For action recognition tasks, this need is reflected in the dynamic nature of video data, which requires models to effectively understand context and temporal relationships. Consequently, existing methods primarily leverage pre-trained CLIP models, enhancing them with carefully designed prompts [37], adapters [44, 25, 41], and temporal modules [24] to enable the model to acquire more domain-specific knowledge about videos. For example, XCLIP [24] introduces cross-frame attention and multi-frame integration modules for temporal modeling. ActionCLIP [34], by utilizing pre-training to obtain powerful representations, applies prompt engineering to bring action recognition tasks closer to the pre-trained model, followed by end-to-end fine-tuning on the target dataset to achieve superior performance. Although the model capacity brought by a large number of learnable parameters enables better fitting of video-specific inductive biases, it may catastrophically forget the original generalization knowledge, thus reducing the generalizability of models.
@@
On the other hand, to enhance the performance of CLIP on video understanding tasks, Wang et al. [36] proposed ViCLIP, which builds upon CLIP and is trained on a large-scale video-text pair dataset. ViCLIP replaces the native attention mechanism in the visual encoder with spatiotemporal attention, while keeping other design components unchanged. However, to date, no work has explored how to leverage ViCLIP for efficient transfer learning, and methods designed for CLIP cannot be directly applied to ViCLIP.
@@
Thus, is there a method that can efficiently fine-tune pre-trained model without damaging its generalizability? A natural idea is to directly apply the efficient transfer methods of image/text foundation models (i.e., LoRA [17] and AdaptFormer [5]) to video tasks. However, these methods are all based on uni-modal information, making it difficult to adapt to multi-modal models. A simple approach is to apply adapters separately to the two modalities. However, this method overlooks the relationship between text and video representations. Therefore, directly applying the same adapters may not be sufficient to capture task-specific cues that vary between vision and language. Furthermore, existing methods do not fully consider the characteristics of text and video representations. Transfer learning often involves a trade-off between discrimination and generalization—specifically, fine-tuning features that are task-specific and discriminative, while preserving those that are more broadly applicable across tasks.
@@
To this end, we propose a novel Mutli-modal Spatio-Temporal Adapter (MSTA) architecture for multi-modal foundation models such that the text and video representations can be better aligned. As shown in Table 1, our method achieves the best generalization performance with a minimal number of parameters. Specifically, our MSTA includes independent projection layers for the text and video branches to learn task-specific knowledge for each modality. To achieve effective alignment between modalities, we introduce a shared unified feature projection layer, which is jointly utilized by both modalities. During the fine-tuning phase, this shared feature space can receive gradients from both modalities [40], thereby optimizing the alignment between them. For the video branch, we further design two simple up-projection layers to enhance the adaptation capabilities for temporal and spatial features. To further mitigate over-fitting and enhance the generalizability, we introduce a spatio-temporal description-guided consistency constraint. This method successfully transfers knowledge from a frozen encoder to a learnable encoder through knowledge distillation, allowing the model to maintain the generalization strength of the pre-trained base model when handling new tasks in few-shot scenarios [30]. Specifically, we impose a consistency constraint on the text branch between the trainable model with MSTA and the pre-trained model. We leverage a pre-trained large language model (e.g., DeepSeek [7]) to generate more detailed and descriptive sentences of temporal and spatial features, and impose a consistency constraint between the representations of the learnable text encoder and the pre-trained text encoder based on these sentences. To validate the effectiveness of our model, we perform extensive experiments across six benchmark datasets: Kinetics-400 [1], Kinetics-600 [2], UCF-101 [31], HMDB-51 [20], SomethingSomething V2 [13], and ActivityNet [15]. The results show that our approach attains state-of-the-art performance in open-vocabulary tasks, such as zero-shot and few-shot learning, and consistently enhances performance when integrated with existing pipelines in fully supervised settings. The key contributions of this work are as follows:
@@ 3.1 Preliminaries
Figure 2: The proposed Multi-modal Spatil-Temporal Adapter (MSTA) for the transformer-based multi-modal foundation models. Only the additional adapters are optimized, while the entire pretrained CLIP model remains frozen. To achieve a balance between the trade-off of discrimination and generalization, we selectively tune only a few higher layers ($\geq k$). MSTA utilizes shared weights across video and text representations, allowing the model to capture common cues from both branches.
@@ 3.2 Multi-modal Spatio-Temporal Adapter
Holistic design. As shown in Fig. 2, different from most of existing methods that add adapters or tokens to the whole network or some lower layers, the new adapter A are partially added into a few higher-layers of both image and text encoders. Formally, for the image encoder V, we add our adapters $A_{v}$ from the $k$-th transformer block and modified Eq. (2) as follows:
As illustrated in Fig. 2, unlike most existing methods that introduce adapters or tokens across the entire network or into lower layers, our adapter $\mathcal{A}$ is selectively incorporated into only a few higher layers of both the video and text encoders. Specifically, for the video encoder $E_{v}$, we incorporate adapters $\{\mathcal{A}_{v}^{j}\}_{j=k}^{L}$ starting from the $k$-th transformer block, modifying Eq. (2) as follows:
@@
Figure 3: Overview of the proposed Multi-modal Spatil-Temporal Adapter (MSTA).
@@
Here, $\mathbf{W}^{ku}$ and $\mathbf{W}^{kd}$ represent the ”Up” and ”Down” projection layers of the $k$-th layer, as illustrated in Fig. 3, with the modality branch indicated by the subscripts. $\mathbf{W}_{v}^{ku-s}$ is the spatial up-projection layer implemented by a linear layer, while $\mathbf{W}_{v}^{ku-t}$ is the temporal up-projection layer implemented by a 3D convolution layer. $\mathbf{W}^{ks}$ denotes the $k$-th projection layer, which is shared across different branches.
@@ 3.3 Spatio-Temporal Description-guided Consistency Constraint
Table 2: Base-to-novel generalization: We compare the generalization ability of MSTA with models that adapt ViCLIP for video tasks. “HM” is the harmonic mean of base and novel accuracy, providing the trade-off between adaption and generalization. ”Zero-ViCLIP” refers to zero-shot ViCLIP, which directly uses the pre-trained ViCLIP for inference.
K-400 HMDB-51 UCF-101 SSv2
Method Base Novel HM Base Novel HM Base Novel HM Base Novel HM
@@
Vanilla CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICML^{\prime}21)}}$ 55.1 55.2 55.1 50.6 48.1 49.3 78.0 63.8 70.2 4.8 5.4 5.1
@@
ActionCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(arXiv^{\prime}22)}}$ 61.5 47.2 53.4 69.3 38.3 49.3 90.5 58.7 71.2 13.5 10.4 11.7
XCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ECCV^{\prime}22)}}$ 74.7 56.7 64.5 69.9 45.3 55.0 90.7 59.5 71.9 8.9 6.8 7.7
A5${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ECCV^{\prime}22)}}$ 69.9 38.4 49.6 46.9 16.5 24.4 90.9 40.8 56.3 8.6 6.4 7.3
@@
Vanilla CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICML^{\prime}21)}}$ 76.4 63.0 69.1 71.5 55.3 62.4 92.3 67.9 78.2 13.2 10.8 11.9
ViFi-CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(CVPR^{\prime}23)}}$ 76.8 61.3 68.2 74.5 53.9 62.5 92.5 68.5 78.7 16.5 12.4 14.2
@@
ViCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICLR^{\prime}23)}}$ 78.5 65.3 71.3 77.9 52.4 62.7 95.1 71.5 81.6 19.4 15.2 17.0
@@
Zero-ViCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICLR^{\prime}23)}}$ 68.7 63.4 65.9 62.5 48.9 54.9 81.7 69.8 75.3 9.3 8.9 9.1
+ AdaptFormer${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}22)}}$ 77.7 65.6 71.1 77.5 54.9 64.3 94.9 72.7 82.3 20.4 14.6 17.0
+ LoRA${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}22)}}$ 77.6 65.3 70.9 77.2 54.6 64.0 95.1 72.3 82.1 19.5 13.6 16.0
+ MSTA 78.2 66.2 71.7 77.2 57.5 65.9 95.6 72.8 82.7 21.9 16.3 18.7
+ MSTA + $\mathcal{L}_{\rm CC}$ 78.5 66.5 72.0 77.5 57.9 66.3 96.0 72.9 82.9 22.2 16.5 18.9
@@
Vanilla CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICML^{\prime}21)}}$ 37.2 37.2 37.2 37.2 62.8 62.8 62.8 62.8 2.8 2.8 2.8 2.8
@@
XCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ECCV^{\prime}22)}}$ 53.5 57.8 62.9 64.5 49.3 76.2 84.1 91.6 4.4 5.2 6.6 11.1
A5${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ECCV^{\prime}22)}}$ 41.2 51.3 56.5 62.7 71.7 80.3 85.9 90.3 4.5 5.3 6.3 9.9
@@
Vanilla CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICML^{\prime}21)}}$ 57.8 61.2 65.6 66.7 81.8 86.3 90.1 92.5 6.7 7.3 9.0 12.2
ViFi-CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(CVPR^{\prime}23)}}$ 57.5 62.9 64.7 66.6 80.5 85.3 90.2 93.2 6.5 7.7 8.3 12.9
@@
MoTE${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}24)}}$ 61.0 63.7 66.9 68.5 88.1 90.7 92.2 93.7 7.4 8.7 9.8 12.5
ViCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICLR^{\prime}23)}}$ 53.7 60.4 64.5 70.3 83.0 88.0 92.1 93.2 8.7 9.7 11.6 15.4
@@
Zero-ViCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICLR^{\prime}23)}}$ 47.8 47.8 47.8 47.8 71.0 71.0 71.0 71.0 5.1 5.1 5.1 5.1
+ AdaptFormer${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}22)}}$ 60.0 61.8 66.6 70.1 84.8 88.7 91.7 93.3 8.5 10.1 11.5 15.2
+ LoRA${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}22)}}$ 59.4 61.4 66.5 69.4 83.5 88.3 90.9 92.7 8.4 9.5 11.1 14.9
+ MSTA 60.1 62.2 66.8 69.9 85.1 89.0 92.3 93.5 9.0 10.1 12.7 16.6
+ MSTA + $\mathcal{L}_{\rm CC}$ 61.2 62.3 67.1 70.4 86.1 90.2 92.7 94.8 9.1 10.4 13.4 17.5

### v4

@@ Abstract
Pre-trained vision-language models provide a robust foundation for efficient transfer learning across various downstream tasks. In the field of video action recognition, mainstream approaches often introduce additional modules to capture temporal information. Although the additional modules increase the capacity of model, enabling it to better capture video-specific inductive biases, existing methods typically introduce a substantial number of new parameters and are prone to catastrophic forgetting of previously acquired generalizable knowledge. In this paper, we propose a parameter-efficient Multi-modal Spatio-Temporal Adapter (MSTA) to enhance the alignment between textual and visual representations, achieving a balance between generalizable knowledge and task-specific adaptation. Furthermore, to mitigate over-fitting and enhance generalizability, we introduce a spatio-temporal description-guided consistency constraint. This constraint involves providing template inputs (e.g., "a video of {cls}") to the trainable language branch and LLM-generated spatio-temporal descriptions to the pre-trained language branch, enforcing output consistency between the branches. This approach reduces overfitting to downstream tasks and enhances the distinguishability of the trainable branch within the spatio-temporal semantic space. We evaluate the effectiveness of our approach across four tasks: zero-shot transfer, few-shot learning, base-to-novel generalization, and fully-supervised learning. Compared to many state-of-the-art methods, our MSTA achieves outstanding performance across all evaluations, while using only 2-7% of the trainable parameters in the original model.
@@ 1 Introduction
Multi-modal foundation models (e.g., CLIP [28], ViCLIP [36], BLIP2 [21]), trained on large-scale paired multi-modal datasets, have demonstrated exceptional generalization capabilities. The semantic visual concepts enabled by extensive multi-modal alignment can be effectively transferred to various downstream tasks, such as few-shot image/video classification [19, 3, 43, 29], open-vocabulary detection [12, 39], and segmentation [23]. In this paper, we focus on transferring the pre-trained multi-modal foundation models model (i.e., CLIP [28] and ViCLIP [36]) to video action recognition tasks, highlighting their significant potential to advance research in this domain.
@@
The key to the adaptation process of foundational multi-modal models lies in injecting domain-specific expertise. In video action recognition, this need is particularly evident due to the dynamic and temporal nature of video data, requiring models to effectively capture context and temporal relationships. Consequently, existing methods typically enhance pre-trained CLIP models by integrating carefully designed prompts [37], adapters [44, 25, 41], or temporal modules [24] to better acquire video-specific knowledge. For instance, XCLIP [24] employs cross-frame attention and multi-frame integration modules to model temporal information. ActionCLIP [34] uses pre-training to generate robust representations, applying prompt engineering to align the action recognition task closely with the pre-trained model, followed by fine-tuning to achieve superior performance. Despite the increased model capacity resulting from numerous learnable parameters, such approaches can lead to catastrophic forgetting, thereby diminishing the generalizability of the original models.
@@
Alternatively, Wang et al. [36] proposed ViCLIP to enhance CLIP performance specifically for video understanding. ViCLIP is trained on extensive video-text pairs and incorporates spatiotemporal attention within the visual encoder, retaining other CLIP components unchanged. However, current research has not yet explored effective transfer learning methods specifically designed for ViCLIP, and existing CLIP-based methods cannot be directly applied.
@@
Thus, the critical question arises: Is there an efficient fine-tuning method for pre-trained models that preserves generalizability? One possible approach is directly applying efficient transfer methods used for image/text foundation models, such as LoRA [17] and AdaptFormer [5], to video tasks. However, these methods primarily rely on uni-modal information, making them unsuitable for multi-modal models. Another simple strategy involves independently applying adapters to each modality, but this overlooks the interrelation between text and video representations. Consequently, directly applying identical adapters may not effectively capture task-specific nuances that vary across vision and language modalities. Furthermore, existing approaches do not fully address the distinct characteristics inherent in text and video representations. Effective transfer learning typically involves balancing discrimination and generalization—fine-tuning task-specific discriminative features while preserving features broadly applicable across different tasks.
@@
To this end, we propose a novel Mutli-modal Spatio-Temporal Adapter (MSTA) architecture for multi-modal foundation models such that the text and video representations can be better aligned. As shown in Fig. 1, our method achieves the best generalization performance with a minimal number of parameters. Specifically, our MSTA includes independent projection layers for the text and video branches to learn task-specific knowledge for each modality. To achieve effective alignment between modalities, we introduce a shared unified feature projection layer, which is jointly utilized by both modalities. During the fine-tuning phase, this shared feature space can receive gradients from both modalities [40], thereby optimizing the alignment between them. For the video branch, we further design two simple up-projection layers to enhance the adaptation capabilities for temporal and spatial features. To further mitigate over-fitting and enhance the generalizability, we introduce a spatio-temporal description-guided consistency constraint. This method successfully transfers knowledge from a frozen encoder to a learnable encoder through knowledge distillation, allowing the model to maintain the generalization strength of the pre-trained base model when handling new tasks in few-shot scenarios [30]. Specifically, we impose a consistency constraint on the text branch between the trainable model with MSTA and the pre-trained model. We leverage a pre-trained large language model (e.g., DeepSeek [7]) to generate more detailed and descriptive sentences of temporal and spatial features, and impose a consistency constraint between the representations of the learnable text encoder and the pre-trained text encoder based on these sentences. To validate the effectiveness of our model, we perform extensive experiments across six benchmark datasets: Kinetics-400 [1], Kinetics-600 [2], UCF-101 [31], HMDB-51 [20], SomethingSomething V2 [13], and ActivityNet [15]. The results show that our approach attains state-of-the-art performance in open-vocabulary tasks, such as zero-shot and few-shot learning, and consistently enhances performance when integrated with existing pipelines in fully supervised settings.
The key contributions of this work are as follows:
@@ 3.1 Preliminaries
Figure 2: The proposed Multi-modal Spatio-Temporal Adapter (MSTA) is designed for transformer-based multi-modal foundation models. It optimizes only the additional adapters, keeping the pretrained CLIP model frozen. To balance discrimination and generalization, we selectively fine-tune a few higher layers ($\geq k$). MSTA employs shared weights for video and text representations, enabling the model to capture shared cues from both modalities.
@@ 3.2 Multi-modal Spatio-Temporal Adapter
Holistic design. As illustrated in Fig. 2, unlike most existing methods that introduce adapters or tokens across the entire network or into lower layers, our adapter $\mathcal{A}$ is selectively incorporated into only a few higher layers of both the video and text encoders. Specifically, for the video encoder $E_{v}$, we incorporate adapters $\{\mathcal{A}_{v}^{j}\}_{j=k}^{L}$ starting from the $k$-th transformer block, modifying Eq. (2) as follows:
@@
Figure 3: Overview of the proposed Multi-modal Spatio-Temporal Adapter (MSTA).
@@
Here, $\mathbf{W}^{ku}$ and $\mathbf{W}^{kd}$ represent the "Up" and "Down" projection layers of the $k$-th layer, as illustrated in Fig. 3, with the modality branch indicated by the subscripts. $\mathbf{W}_{v}^{ku-s}$ is the spatial up-projection layer implemented by a linear layer, while $\mathbf{W}_{v}^{ku-t}$ is the temporal up-projection layer implemented by a 3D convolution layer. $\mathbf{W}^{ks}$ denotes the $k$-th projection layer, which is shared across different branches.
@@ 3.3 Spatio-Temporal Description-guided Consistency Constraint
Table 2: Base-to-novel generalization: We compare the generalization ability of MSTA with models that adapt ViCLIP for video tasks. “HM” is the harmonic mean of base and novel accuracy, providing the trade-off between adaption and generalization. "Zero-ViCLIP" refers to zero-shot ViCLIP, which directly uses the pre-trained ViCLIP for inference.
Method GFLOPs K-400 HMDB-51 UCF-101 SSv2
Base Novel HM Base Novel HM Base Novel HM Base Novel HM
@@
Vanilla CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICML^{\prime}21)}}$ 281 55.1 55.2 55.1 50.6 48.1 49.3 78.0 63.8 70.2 4.8 5.4 5.1
@@
ActionCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(arXiv^{\prime}22)}}$ 282 61.5 47.2 53.4 69.3 38.3 49.3 90.5 58.7 71.2 13.5 10.4 11.7
XCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ECCV^{\prime}22)}}$ 145 74.7 56.7 64.5 69.9 45.3 55.0 90.7 59.5 71.9 8.9 6.8 7.7
A5${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ECCV^{\prime}22)}}$ 284 69.9 38.4 49.6 46.9 16.5 24.4 90.9 40.8 56.3 8.6 6.4 7.3
@@
Vanilla CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICML^{\prime}21)}}$ 281 76.4 63.0 69.1 71.5 55.3 62.4 92.3 67.9 78.2 13.2 10.8 11.9
ViFi-CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(CVPR^{\prime}23)}}$ 281 76.8 61.3 68.2 74.5 53.9 62.5 92.5 68.5 78.7 16.5 12.4 14.2
@@
ViCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICLR^{\prime}23)}}$ 161 78.5 65.3 71.3 77.9 52.4 62.7 95.1 71.5 81.6 19.4 15.2 17.0
@@
Zero-ViCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICLR^{\prime}23)}}$ 161.74 68.7 63.4 65.9 62.5 48.9 54.9 81.7 69.8 75.3 9.3 8.9 9.1
+ AdaptFormer${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}22)}}$ 179.80 77.7 65.6 71.1 77.5 54.9 64.3 94.9 72.7 82.3 20.4 14.6 17.0
+ LoRA${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}22)}}$ 179.80 77.6 65.3 70.9 77.2 54.6 64.0 95.1 72.3 82.1 19.5 13.6 16.0
+ MSTA 179.84 78.2 66.2 71.7 77.2 57.5 65.9 95.6 72.8 82.7 21.9 16.3 18.7
+ MSTA + $\mathcal{L}_{\rm CC}$ 179.84 78.5 66.5 72.0 77.5 57.9 66.3 96.0 72.9 82.9 22.2 16.5 18.9
@@
Vanilla CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICML^{\prime}21)}}$ 281 37.2 37.2 37.2 37.2 62.8 62.8 62.8 62.8 2.8 2.8 2.8 2.8
@@
XCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ECCV^{\prime}22)}}$ 145 53.5 57.8 62.9 64.5 49.3 76.2 84.1 91.6 4.4 5.2 6.6 11.1
A5${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ECCV^{\prime}22)}}$ 284 41.2 51.3 56.5 62.7 71.7 80.3 85.9 90.3 4.5 5.3 6.3 9.9
@@
Vanilla CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICML^{\prime}21)}}$ 282 57.8 61.2 65.6 66.7 81.8 86.3 90.1 92.5 6.7 7.3 9.0 12.2
ViFi-CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(CVPR^{\prime}23)}}$ 281 57.5 62.9 64.7 66.6 80.5 85.3 90.2 93.2 6.5 7.7 8.3 12.9
@@
MoTE${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}24)}}$ 141 61.0 63.7 66.9 68.5 88.1 90.7 92.2 93.7 7.4 8.7 9.8 12.5
ViCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICLR^{\prime}23)}}$ 161 53.7 60.4 64.5 70.3 83.0 88.0 92.1 93.2 8.7 9.7 11.6 15.4
@@
Zero-ViCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICLR^{\prime}23)}}$ 161 47.8 47.8 47.8 47.8 71.0 71.0 71.0 71.0 5.1 5.1 5.1 5.1
+ AdaptFormer${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}22)}}$ 179.80 60.0 61.8 66.6 70.1 84.8 88.7 91.7 93.3 8.5 10.1 11.5 15.2
+ LoRA${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}22)}}$ 179.80 59.4 61.4 66.5 69.4 83.5 88.3 90.9 92.7 8.4 9.5 11.1 14.9
+ MSTA 179.84 60.1 62.2 66.8 69.9 85.1 89.0 92.3 93.5 9.0 10.1 12.7 16.6
+ MSTA + $\mathcal{L}_{\rm CC}$ 179.84 61.2 62.3 67.1 70.4 86.1 90.2 92.7 94.8 9.1 10.4 13.4 17.5